# WaveGNN: Modeling Irregular Multivariate Time Series for Accurate Predictions

### v1

@@ Abstract
Accurately modeling and analyzing time series data is crucial for downstream applications across various fields, including healthcare, finance, astronomy, and epidemiology. However, real-world time series often exhibit irregularities such as misaligned timestamps, missing entries, and variable sampling rates, complicating their analysis. Existing approaches often rely on imputation, which can introduce biases. A few approaches that directly model irregularity tend to focus exclusively on either capturing intra-series patterns or inter-series relationships, missing the benefits of integrating both. To this end, we present WaveGNN, a novel framework designed to directly (i.e., no imputation) embed irregularly sampled multivariate time series data for accurate predictions. WaveGNN utilizes a Transformer-based encoder to capture intra-series patterns by directly encoding the temporal dynamics of each time series. To capture inter-series relationships, WaveGNN uses a dynamic graph neural network model, where each node represents a sensor, and the edges capture the long- and short-term relationships between them. Our experimental results on real-world healthcare datasets demonstrate that WaveGNN consistently outperforms existing state-of-the-art methods, with an average relative improvement of $14.7\%$ in F1-score when compared to the second-best baseline in cases with extreme sparsity. Our ablation studies reveal that both intra-series and inter-series modeling significantly contribute to this notable improvement.
1 Introduction
@@
The widespread use of sensors across various sectors, including healthcare, finance, astronomy, and urban applications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], has led to a significant increase in the collection of time series data, ranging from univariate to multivariate types, where multiple measures are recorded simultaneously. Effectively modeling and analyzing these time series datasets is crucial for downstream applications such as monitoring patient health, optimizing treatments, and managing disease outbreaks, where accurate and timely data interpretation can significantly impact outcomes. In real-world applications, though, time series often exhibit irregularities, frequently caused by sensor malfunctions, different sampling rates across sensors, and cost-saving measures [19]. This variability in data collection renders traditional machine learning (ML) approaches, which typically rely on a fixed number of regularly sampled observations across all sensors, ineffective for accurate predictions [20, 19].
Modeling irregularly sampled time series can be complex due to several factors: measurements from different sensors may not align, data may be missing, and each measurement sequence may vary in size [19, 21, 22]. To address these challenges, recent developments have concentrated on creating specialized algorithms and model architectures designed to effectively manage these irregularities [23].
One category of approaches tackles irregularity by filling in the missing values, thereby converting the datasets into regularly sampled time series. These imputation techniques vary from traditional hand-crafted imputation methods [24] to more advanced learned interpolation methods [25, 21]. Once the missing values are filled, the resulting regular time series are utilized to perform the downstream predictions [26, 27, 28]. While these techniques address data gaps, they introduce biases and distortions in the data distribution. Furthermore, the absence of measurements can itself provide valuable insights; for instance, periodically missing heart rate data might indicate that a patient underwent a medical procedure during which the monitoring device was removed.
Another category of approaches focuses on directly learning from the irregularities by designing specialized model architectures [21, 29, 19, 20]. For instance, the GRU-D model incorporates a decaying mechanism within gated recurrent units (GRUs) to better handle irregular sampling intervals [29], while mTAND utilizes multi-time attention to analyze non-uniform measurements effectively [21]. These techniques predominantly capture the correlations within a single sensor’s measurements over time, focusing on intra-series dynamics. Meanwhile, other methods have been proposed to capture inter-series dependencies, modeling explicitly the relationships across sensors within the dataset [19].
Despite these advancements, most approaches tend to specialize in either capturing intra-series dynamics or inter-series relationships, but rarely both simultaneously. Yet, effectively modeling both types of relationships must lead to more accurate predictions and deeper insights. For example, understanding both the individual progression of patients’ vital signs (e.g., heart rate) and how they interact with one another (e.g., the co-dependencies between heart rate and blood pressure over time) can provide a holistic view of a patient’s condition, potentially improving diagnostic accuracy and treatment outcomes. Additionally, in scenarios where one or more sensor observations are missing, irregularity can be compensated by exploiting insights among the correlated sensors.
Inspired by these insights, we introduce WaveGNN, a novel framework designed to directly (without imputation) embed irregularly sampled multivariate time series datasets for accurate predictions. WaveGNN captures both the intra-series patterns and the inter-series relationships simultaneously, with two main components. To effectively capture the temporal dependencies of each individual time series, the first component is a Transformer-based encoder. This encoder incorporates a masked attention mechanism that excludes missing data observations during processing, along with a relative timestamp encoding mechanism to capture the time intervals between consecutive observations. Given that more recent observations can be more indicative of the final prediction, a decay mechanism is also incorporated within the encoder to prioritize newer observations. To adequately capture inter-sensor dependencies, the second component of WaveGNN is a dynamic graph neural network, where each node corresponds to a sensor in the dataset, and the edges between nodes denote the correlations between these sensors. These edges are dynamically inferred based on sequence embeddings generated by the transformer-based encoder during the time window and learned global node embeddings shared across all samples and all time windows. Consequently, WaveGNN effectively captures both the dynamic short-term relationships observed within each window and the long-term relationships across the entire dataset, effectively mitigating the impact of missing data by leveraging inter-sensor information.
We evaluate WaveGNN across four real-world healthcare datasets, each associated with distinct classification tasks. Our experiments show that WaveGNN consistently outperforms both imputation-based and specialized state-of-the-art baselines by up to 10.2% in AUPRC and 3.4% in accuracy. Even in real-world scenarios where entire sets of sensor observations are missing, WaveGNN maintains high performance, outperforming baselines with an average relative improvement of 14.7% in F1 score on the challenging PAM dataset, which involves an 8-class classification task. Our ablation study reveals that each proposed component contributes to the overall effectiveness of WaveGNN, with inter-series and intra-series dependencies showing the most significant performance drops when excluded, leading to F1 score reductions of 33.8% and 13.2%, respectively.
2 Related Work
Approaches for handling irregularities in multivariate time series vary widely. Temporal Discretization-based Embedding (TDE) methods transform irregular time series into regular forms through imputation [26]. Such approaches enable the use of standard deep learning models but inherently introduce biases and distort data distributions [27, 28, 24]. To address their shortcomings, approaches such as IP-Net and DGM2 proposed using kernel functions to interpolate data against reference points to achieve temporal alignment [25, 30]. Shukla et al. further proposed a time attention mechanism with time embeddings to learn interpolation representations [21], while Zhang et. al [26] introduced a hybrid model that combines different TDE methods and integrated hand-crafted imputation embeddings into learned interpolation embeddings to improve medical predictions [26]. These approaches still fundamentally rely on imputation, retaining the inherent drawbacks of TDE methods.
Another group of approaches model irregularity directly through specialized algorithms or model architecture designs. To that end, GRU-D employs decaying mechanisms within GRUs to handle the irregular sampling of the data [29]. SeFT applies set function learning to time series classification, treating the series as a set of observations rather than a sequence [20]. Additionally, Raindrop employs graph neural networks to conceptualize multivariate time series as separate sensor graphs, capturing complex inter-variable relationships [19]. While these approaches avoid the TDE drawbacks, they often require extensive customization and typically focus on capturing either intra-variable dynamics or inter-variable dependencies, but not both, potentially overlooking essential data patterns. In contrast, WaveGNN introduces a more integrated approach, utilizing a standard transformer architecture with masking and time encoding to effectively model intra-series relationships and a graph neural network architecture to capture short and long-term inter-variable dependencies simultaneously. We evaluate WaveGNN against the aforementioned approaches in Section 5.2.
Graph-based approaches have been extensively applied to model multivariate time series datasets. As an example of dynamic graph-based approaches, BysGNN creates a dynamic multi-context graph for forecasting visits to points of interest (POIs) based on spatial, temporal, semantic, and taxonomic contexts [6]. Similarly, NeuroGNN, captures multi-context correlations in EEG data for seizure classification [4]. These approaches are designed for regular time series data and rely on having external context information to accurately infer inter-series relationships. In contrast, WaveGNN models the inter-series relationships given irregular multivariate time series without requiring additional context.
Figure 1: Overall WaveGNN Framework
3 Preliminaries
@@
A time series observation, $s^{t}_{i,v}$, refers to a single measurement recorded by sensor $v$ at timestamp $t$ for sample $i$. Sequentially recorded observations over time form a time series $s_{i,v}$ for that sample and sensor.
@@
Problem Definition (Irregular Multivariate Time Series Prediction Task) Given a dataset $D=\{(S_{i},t_{i},p_{i},y_{i})\}_{i=1}^{N}$ of irregularly sampled multivariate time series, the goal is to learn a function $f:S_{i}\rightarrow g_{i}$ that maps the sample $s_{i}$, to a fixed-length representation $g_{i}$. This representation can be utilized to predict a label $\hat{y_{i}}\in\{1,\ldots,C\}$ relevant to the downstream task.
4 Methodology
Figure 1 presents the end-to-end pipeline of WaveGNN. In the first step, WaveGNN allocates nodes for each sensor, initializing their states using static features from the dataset, like patient demographics, if available, or randomly if not. Next, each sensor’s irregular observation sequence is embedded using a customized Transformer encoder. This encoder employs masked attention, modified temporal encoding, and a decay mechanism to handle irregularities and capture correlations within each sensor’s sequence. The node states are updated based on these sequence embeddings to represent the latent state of each sensor after the observed window. In the third step, WaveGNN establishes edges that represent the correlations across sensors to effectively utilize inter-series correlations. This allows WaveGNN to compensate for missing observations during the input window by leveraging information from other related sensors to each sensor. These edges are formed based on the updated node states, capturing short-term relationships during the observed window, and on learned global node embeddings, capturing long-term relationships derived from the entire training dataset. Next, WaveGNN performs message passing, which updates the node states to incorporate information from these inferred relationships. This process allows each sensor to benefit from the data captured by other sensors, reducing the impact of irregularity. Finally, the combined node states are aggregated into a single graph-level embedding vector, which is then fed into the prediction head to produce the final prediction. The detailed steps are presented in the following.
4.1 Sensor Graph Construction
Initialization. In WaveGNN, the graph initialization step establishes the preliminary graph structure without connecting any nodes. These will later be updated to reflect the relationships and state of the given sample during the observed time window. The graph comprises nodes, each for a different sensor involved in measuring a time-dependent feature.
For each sample $i$, each node is initialized with a state vector derived from static features $p_{i}$ corresponding to the sample (e.g., demographics such as age and gender and clinical contexts such as ICU types in the medical downstream tasks). If this information is not available for a task, the vectors are initialized randomly. These features are transformed using a two-layer MLP with ReLU activations to create high-dimensional representations that encapsulate the sample’s baseline characteristics and provide a sample-specific initial context for the task:
@@ 4.1 Sensor Graph Construction
$\bullet$ Static Long-Term Similarities ($A_{L}$): WaveGNN learns global node embeddings $\tilde{Z}=\{\tilde{z}_{1},\dots,\tilde{z}_{n}\}\in\mathbb{R}^{n\times M}$, where $M$ is the embedding dimension. These embeddings reflect long-term similarities between sensors, learned from the entire dataset for a specific downstream task. The $A_{L}$ matrix is then computed using the dot product between these global node embeddings, followed by a softmax normalization:
@@
4.2 Graph-based Prediction
With the updated node states $Z$ and adjacency matrix $A$, the graph $G=(Z,A)$ for the given window of observations is constructed, encapsulating the latest node states and their inter-relationships. We perform GNN-based message passing on this graph to capture the node relationships within the node embeddings. Afterward, we apply pooling and transformations to obtain a single embedding vector representing all important node embeddings. The final embedding is used to make the prediction.
@@ 4.2 Graph-based Prediction
Graph Embedding. To derive a graph-level representation from the node embeddings, two operations are performed: $\bullet$ Max Pooling: Applied to the node embeddings to extract the most significant features, helping to capture the most expressive node attributes relevant to the prediction task.
@@
5 Experiments
5.1 Experimental Setup
Datasets and Evaluation Setup. We evaluate WaveGNN on four real-world healthcare datasets. P12 [37] and MIMIC III [38] contain ICU patient data, with labels indicating survival during hospitalization (48-IHM). MIMIC III also includes labels for multilabel phenotype classification (24-PHE). P19 [39] focuses on sepsis detection, while PAM [40] is an activity monitoring dataset with labels for different activities. P12, P19, and MIMIC III are naturally irregular datasets, while irregularity in PAM is introduced artificially. The dataset statistics’ are presented in Table 1. For a fair comparison, we followed the same preprocessing steps with previous studies [19, 28], and adopted the same dataset splits. To that extent, the training, validation, and testing ratios for P12, P19, and PAM is 80:10:10, and 85:7.5:7.5 for MIMIC III. The indices of these splits are fixed across all methods and experiments. Evaluation metrics also match those used in prior studies to maintain consistency, and all hyperparameters were kept the same across all datasets and experiments. More details regarding the evaluation setup and hyperparameter tuning can be found in the Supplementary Material sections A.1-A.3.
Table 1: Statistics of the datasets used in experimental study.

### v2

# WaveGNN: Integrating Graph Neural Networks and Transformers for Decay-Aware Classification of Irregular Clinical Time-Series

@@ Abstract
Clinical time series are often irregularly sampled, with varying sensor frequencies, missing observations, and misaligned timestamps. Prior approaches typically address these irregularities by interpolating data into regular sequences, thereby introducing bias, or by generating inconsistent and uninterpretable relationships across sensor measurements, complicating the accurate learning of both intra-series and inter-series dependencies. We introduce WaveGNN, a model that operates directly on irregular multivariate time series without interpolation or conversion to a regular representation. WaveGNN combines a decay-aware Transformer to capture intra-series dynamics with a sample-specific graph neural network that models both short-term and long-term inter-sensor relationships. Therefore, it generates a single, sparse, and interpretable graph per sample. Across multiple benchmark datasets (P12, P19, MIMIC-III, and PAM), WaveGNN delivers consistently strong performance, whereas other state-of-the-art baselines tend to perform well on some datasets or tasks but poorly on others. While WaveGNN does not necessarily surpass every method in every case, its consistency and robustness across diverse settings set it apart. Moreover, the learned graphs align well with known physiological structures, enhancing interpretability and supporting clinical decision-making.
I Introduction
@@
Continuous monitoring devices, electronic health‐record (EHR) systems, and wearable sensors now produce vast clinical time-series. Effectively modeling these datasets is crucial for downstream applications such as precision medicine, monitoring patient health, optimizing treatments, and managing disease outbreaks. Unlike the ideal, uniformly sampled sequences assumed in classical time-series analysis, clinical measurements often exhibit irregularities, caused by sensor malfunctions, different sampling rates across sensors, and cost-saving measures [1]. This variability in data collection renders traditional machine learning approaches, which rely on a fixed number of regularly sampled observations across all sensors, ineffective for accurate predictions [2, 1].
Understanding clinical time series requires capturing both the individual progression of each physiological signal and how different signals interact over time. For example, tracking heart rate in isolation provides only partial information, whereas modeling its temporal relationship with blood pressure can reveal important clinical patterns. Incorporating such inter-series dependencies is essential for improving diagnostic accuracy. Consequently, modelling these data poses two coupled challenges: (1) Intra-series irregularity, where each sensor records observations at uneven intervals and frequently experiences missing measurements. (2) Inter-series discrepancy, where different sensors operate at varying frequencies, resulting in misaligned, unequal-length sequences. These complexities significantly hinder traditional modeling methods, emphasizing the need for novel approaches capable of effectively addressing both types of irregularities.
A large body of prior work addresses intra-series irregularity. One line of methods applies interpolation or imputation to fill missing values, converting irregular series into regularly sampled sequences for downstream models [3, 4, 5]. While effective in some cases, these methods risk distorting temporal patterns and removing information carried by the structure of missingness itself. Alternative approaches such as GRU-D [6] and Neural ODEs [7] avoid imputation by directly modeling irregular time intervals, but they typically operate on each sensor independently and do not account for inter-series interactions. This independence overlooks the fact that measurements from related sensors often exhibit strong dependencies (e.g., correlations between heart rate and respiration), so leveraging inter-series dynamics can help impute missing values more accurately and preserve meaningful temporal patterns. Without modeling such inter-series relationships, intra-series imputation risks being less reliable, particularly when data sparsity is severe.
Recent work has attempted to model inter-variable dependencies in irregular time series, but these approaches introduce key limitations. Many methods, such as UTDE [8], T-PatchGNN [9], and Warpformer [10], bypass irregularity by first converting the entire irregular MTS into a regularly spaced representation. UTDE applies learned interpolation to fill in missing values, T-PatchGNN divides each signal into fixed-length patches, and Warpformer transforms signals onto a shared grid by downsampling high-frequency sensors and upsampling sparse ones. These transformations can distort temporal patterns and remove information carried by missingness. They also introduce additional hyperparameters such as patch size, stride, and grid resolution, which require tuning for each dataset. Moreover, these models do not produce a consistent or interpretable structure across variables, making it difficult to understand or trust the relationships they learn in critical domains like healthcare.
On the other hand, graphs have long been used to model structured and interpretable relationships. Recently, graph-based architectures have been adopted to explicitly model inter-variable dependencies in irregular time-series, aiming to address inter-series discrepancy. This is particularly desirable in healthcare, where understanding interactions between physiological signals is critical for decision-making [1]. Recent graph-based techniques for modeling clinical time series, T-PatchGNN [9] and GraFITi [11], often face critical limitations: they either construct multiple fragmented graphs that are difficult to aggregate into a coherent representation, or they restrict interactions between variables only through simultaneous observations, failing to model relationships effectively when sensor data are not aligned. Additionally, many of these methods do not produce an interpretable, unified graph structure, undermining transparency and limiting their clinical applicability. We compare with both models in our experiments and analyze their limitations in Section V, alongside a detailed discussion in Section II.
To address the limitations of existing approaches, we propose WaveGNN, a model that jointly captures intra-series irregularity and inter-series dependencies without converting the data to a regular format or introducing task-specific architectural hyperparameters [^1] Code and data pointers are available at https://github.com/USC-InfoLab/WaveGNN/.. WaveGNN combines a Transformer-based temporal encoder with a single dynamic graph neural network that operates over sensors, enabling direct modeling of both intra-series temporal patterns and inter-sensor relationships.
The Transformer-based encoder models each sensor independently. It uses masked attention to ignore missing values during computation and applies relative time encodings to capture the time gaps between observations. To account for the fact that recent measurements are often more relevant than older ones in clinical settings, the encoder includes a learnable decay mechanism that assigns lower weights to older inputs. This design allows the model to process irregular input sequences directly, without requiring imputation, interpolation, or resampling.
On top of the per-sensor representations, the dynamic graph module constructs a sparse, interpretable graph for each sample. Each sensor is treated as a node, and edges are formed based on two signals: (1) local similarity derived from Transformer representations within a time window and (2) global relationships encoded via learnable embeddings shared across all samples. A gating mechanism balances these short-term and long-term signals. The resulting graph reflects both sample-specific and population-level dependencies, and its structure is adapted dynamically for each sample. GNN layers propagate information across this graph to form a compact representation used for downstream classifications. Unlike prior models, WaveGNN produces a single, stable interaction graph per sample that is both clinically interpretable and robust to irregular sampling.
We evaluate WaveGNN on four real-world healthcare datasets covering binary, multi-class, and multi-label classification tasks. Across all benchmarks, WaveGNN demonstrates consistent performance, achieving the lowest average rank (1.67) across 12 evaluations. The model is also highly robust to missing data: across experiments simulating different levels of missingness, WaveGNN outperforms all baselines in 20 out of 24 experiments and ranks second in the remaining four, with relative F1 improvements of up to 12.3% on PAM. Ablation studies further show the necessity of jointly modeling inter- and intra-series dependencies, with F1 drops of 35.3% and 15.1% when these components are removed. Finally, WaveGNN produces interpretable graphs whose learned dependencies align with known physiological relationships (Section V-D), highlighting its utility for clinical decisions.
The remainder of the paper is organized as follows. Section II reviews related work. Section III covers preliminaries and problem definition. Section IV describes the WaveGNN architecture. Section V presents experiments on four clinical datasets and analyzes model interpretability. Section VI concludes the paper.
II Related Work
Modeling Intra-Series Irregularity Intra-series irregularity arises from variable sampling rates, missing values, and uneven observation intervals within each sensor. One line of work addresses this by converting irregular sequences into regularly spaced time series via interpolation. For instance, IP-Net [4] and DGM2-O [12] apply kernel interpolation over fixed reference points, while mTAND [5] combines multi-head interpolation with time-aware attention. These methods support standard architectures but risk introducing bias and losing information inherent in the pattern of missingness [8, 13]. An alternative direction aims to model irregular sequences directly. GRU-D [6] incorporates exponential decay terms into both inputs and hidden states to reflect temporal recency. SeFT [2] uses set functions to generate permutation-invariant embeddings from unordered observations. Neural ODEs [7] model continuous-time dynamics to naturally handle irregular sampling. While these approaches preserve temporal structure, they typically process each sensor independently and do not model cross-variable relationships.
Modeling Inter-Series Dependencies Capturing dependencies across variables is critical in clinical time series, where interactions, such as between heart rate and blood pressure, can indicate physiological states. Several recent methods address this by building inter-variable representations on top of regularized input representation. UTDE [8] jointly learns interpolation functions and applies cross-variable attention within a Transformer. Warpformer [10] resamples signals onto a shared temporal grid via learnable downsampling and upsampling. T-PatchGNN [9] segments each signal into fixed-length regularly-spaced patches and constructs local graphs to capture short-term inter-series dependencies. However, these models rely on regularization that may distort temporal cues and introduce tuning overhead via additional hyperparameters (e.g., patch sizes, strides, or grid resolutions). Moreover, they do not yield consistent or interpretable inter-variable relationship structures.
Graph-Based Approaches Graph-based models offer a natural way to encode structured and interpretable dependencies in multivariate time series. Prior works that provide interpretable relationship structures include BysGNN [14], which constructs context-aware dynamic graphs for spatiotemporal forecasting, and NeuroGNN [15], which captures multi-channel EEG interactions for seizure detection. However, these models are designed for regularly sampled inputs.
To extend graph modeling to irregular settings, Raindrop [1] introduces event-based message passing over variable nodes. However, it lacks explicit intra-series modeling, treats all observations equally, and does not encode long-term dependencies between variables. T-PatchGNN [9] builds patch-specific graphs, but these vary across time and do not produce a coherent interaction graph per instance. GraFITi [11] constructs a bipartite graph over time and variable nodes, but restricts interactions to co-observed timestamps, limiting its effectiveness under heavy misalignment.
In contrast, WaveGNN integrates a decay-aware Transformer for intra-series encoding with a dynamic, per-sample graph over variables that captures both short-term temporal similarity and long-range structural dependencies. This yields a consistent, sparse, and interpretable interaction graph for each instance, without relying on interpolation, handcrafted temporal alignment, or task-specific hyperparameters. Evaluation is presented in Section V.
III Preliminaries
@@
A time series observation, $s^{t}_{i,v}$, is a single measurement recorded by sensor $v$ at timestamp $t$ for sample $i$. Sequential observations over time form a time series $s_{i,v}$ for that sample and sensor.
@@
[…]
Figure 1: Overall WaveGNN Framework
IV Methodology
Figure 1 presents the end-to-end pipeline of WaveGNN. In the first step, WaveGNN allocates nodes for each sensor, initializing their states using static features from the dataset, like patient demographics, if available, or randomly if not. Next, each sensor’s irregular observation sequence is embedded using a customized Transformer encoder. This encoder employs masked attention, modified temporal encoding, and a decay mechanism to handle irregularities and capture correlations within each sensor’s sequence. The node states are updated based on these sequence embeddings to represent the latent state of each sensor after the observed window. In the third step, WaveGNN establishes edges that represent the correlations across sensors to effectively utilize inter-series correlations. This allows WaveGNN to compensate for missing observations during the input window by leveraging information from other related sensors to each sensor. These edges are formed based on the updated node states, capturing short-term relationships during the observed window, and on learned global node embeddings, capturing long-term relationships derived from the entire training dataset. Next, WaveGNN performs message passing, updating node states to incorporate information from inferred relationships. This allows each sensor to benefit from other sensors’ data, reducing the impact of irregularity. Finally, the updated node states are aggregated into a graph-level embedding vector, which is fed into the prediction head. The detailed steps are presented below.
IV-A Sensor Graph Construction
Initialization. In WaveGNN, the graph initialization step establishes the preliminary graph structure without connecting nodes. These will later be updated to reflect the relationships and state of the given sample during the observed time window. The graph comprises nodes, each representing a different sensor measuring a time-dependent feature.
For each sample $i$, each node is initialized with a state vector derived from static features $p_{i}$ corresponding to the sample (e.g., demographics such as age and gender, and clinical contexts such as ICU types in medical downstream tasks). If unavailable, the vectors are initialized randomly. These features are transformed using a two-layer MLP with ReLU activations to create high-dimensional representations that encapsulate the sample’s baseline characteristics and provide a sample-specific initial context for the task:
@@ IV-A Sensor Graph Construction
$\bullet$ Static Long-Term Similarities ($A_{L}$): WaveGNN learns global node embeddings $\tilde{Z}=\{\tilde{z}_{1},\dots,\tilde{z}_{n}\}\in\mathbb{R}^{n\times M}$, where $M$ is the embedding dimension. These embeddings reflect long-term similarities between sensors, learned from the entire dataset for a specific downstream task. $A_{L}$ matrix is then computed using the dot product between these global node embeddings, followed by a softmax normalization:
@@
IV-B Graph-based Prediction
With the updated node states $Z$ and adjacency matrix $A$, the graph $G=(Z,A)$ for the given window of observations is constructed, encapsulating the latest node states and their inter-relationships. We perform GNN-based message passing on this graph to capture the node relationships through the node embeddings. Afterward, we apply pooling and transformations to obtain a single embedding vector representing all important node embeddings. The final embedding is used to make the prediction.
@@ IV-B Graph-based Prediction
Graph Embedding. To derive a graph-level representation from the node embeddings, two operations are performed:
$\bullet$ Max Pooling: Applied to the node embeddings to extract the most significant features, helping to capture the most expressive node attributes relevant to the prediction task.
@@
V Experiments
V-A Experimental Setup
Datasets and Evaluation Setup. We evaluate WaveGNN on four real-world healthcare datasets. P12 [22] and MIMIC III [23] contain ICU patient data, with labels indicating survival during hospitalization (48-IHM). MIMIC III also includes labels for multilabel phenotype classification (24-PHE). P19 [24] focuses on sepsis detection, while PAM [25] is an activity monitoring dataset with labels for different activities. P12, P19, and MIMIC III are naturally irregular datasets, while irregularity in PAM is introduced artificially. The dataset statistics’ are presented in Table I. For a fair comparison, we followed the same preprocessing steps with previous studies [1, 26], and adopted the same dataset splits. To that extent, the training, validation, and testing ratios for P12, P19, and PAM is 80:10:10, and 85:7.5:7.5 for MIMIC III. The indices of these splits are fixed across all methods and experiments. Evaluation metrics also match those used in prior studies to maintain consistency, and all hyperparameters were kept the same across all datasets and experiments.
TABLE I: Statistics of the datasets used in experimental study.